40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 28 23:33:38 2020, gautam, Update, BHD, Some more hardware changes 
    Reply  Sat Aug 29 22:46:29 2020, gautam, Update, BHD, New homodyne-phase control electronics zetaDrive.pdftrekTFs.pdf
       Reply  Wed Sep 2 00:49:47 2020, gautam, Update, BHD, Some notes about homodyne phase heterodyneMICH.pdfunwrappedPhase.pdfunwrappedPhase_zoom.pdfphaseNoisePSD.pdf
          Reply  Thu Sep 3 15:55:04 2020, gautam, Update, BHD, Phase drift between LO and IFO after fiber replacement LOphaseDrift.pdfphaseDrift_tempCorr.png
Message ID: 15545     Entry time: Fri Aug 28 23:33:38 2020     Reply to this: 15549
Author: gautam 
Type: Update 
Category: BHD 
Subject: Some more hardware changes 

Just a quick set of notes detailing changes so that there are no surprises, more details to follow.

  1. Trek driver has been temporarily placed on top of the KEPCO supply east of the OMC electronics rack. Cabling to it has been laid out as well. I turned both off so neither should be energized now.
  2. A new AI chassis (and associated cabling including the DAC SCSI cable and +/-24 V DC cable) has been installed in 1X2.
  3. To map the DAC range to what the Trek driver wants, I've configured the inverting summing amplifier with gain of 1/8. The offset voltage is set to 5V DC instead of 10V as intended, because the DAC can only drive +/-5 V when connected to a single ended receiving/sending unit.
  4. The LO delivery fiber was re-laid, and the interference between the IFO AS beam and LO beams were restored.

I briefly tried some LO PZT mirror dithering tonight, but didn't see the signal. Needs more troubleshooting.

ELOG V3.1.3-